OpenStudy (anonymous):

how do you work this out? 2[(x+7)-9]=5(x-4)

OpenStudy (anonymous):

\[2[x+7-9]=5(x+4)\] \[2[x-2]=5x+20\] \[2x-4=5x+20\] \[2x-5x=20+4\] \[-3x=24\] \[x=24/-3\] \[x=-8\]

OpenStudy (anonymous):

2 (x-2) = 5 (x-4) Expand out terms on both sides: 2 x-4 = 5 x-20 Add (4-5 x) to both sides: -3 x = -16 Divide both sides by -3: x = 16/3